OpenStudy (anonymous):

lim (x^6 + 1,000x^3 + 2,000,000)/(4x6 + 9,000x^3) x→+∞ Calculate the limit algebraically. If the limit does not exist, say why. A)The limit does exist. B)The limit from the left is +∞ whereas the limit from the right is −∞. Since the left and right limits disagree, the limit does not exist. C)The limit from the left is −∞ whereas the limit from the right is +∞. Since the left and right limits disagree, the limit does not exist. D)It is not possible to evaluate a function at x = ∞; therefore the limit does not exist.

OpenStudy (luigi0210):

Have you learned end behavior? Or do you remember it?

OpenStudy (anonymous):

i know when x→+∞ its +∞ but it says its wrong

OpenStudy (anonymous):

and chose D as an answering

OpenStudy (luigi0210):

Try dividing by x^6: \[\LARGE \frac{(x^6 + 1,000x^3 + 2,000,000)}{(4x6 + 9,000x^3)}\]

OpenStudy (luigi0210):

It's hard to do it with latex.. but when you do, you'll get something like: (1+{1000/x^3}+{2000000/x^6})/4+{9000/x^3})

OpenStudy (luigi0210):

Once you do that, plug in infinity: Leaving: \[\LARGE \frac{1}{4}\] any number over infinity is 0.

OpenStudy (luigi0210):

Since infinity is limitless, it will continue to grow, while the numerator stays the same.. eventually the denominator will be so great that it will practically be 0.

OpenStudy (luigi0210):

Also, with end behavior, learned in Algebra II, you could take the highest powers: \[\LARGE \frac{(x^6 + 1,000x^3 + 2,000,000)}{(4x6 + 9,000x^3)}=\frac{x^6}{4x^6} \] Get rid of the x's \[\LARGE \frac{1}{4}\]

OpenStudy (luigi0210):

Okay, I gtg now, hope I made it understandable, good luck!

OpenStudy (anonymous):

thank you